# Ephemeral Rollups are All you Need

### v3

@@ Abstract
In the realm of open and composable gaming, we envision platforms where users actively expand, create, engage, and immerse themselves in a rich world of entertainment. One promising avenue for achieving this vision is through fully on-chain (FOC) games, where both game state and logic reside on the blockchain, maximizing composability. However, we must grapple with inherent limitations and trade-offs, particularly in terms of costs and scalability. This paper proposes a framework that leverages the Solana Virtual Machine (SVM) to scale FOC games without state fragmentation or compromised trust assumptions. The framework introduces a systematic approach for discovering, utilizing, and publishing modular pieces of logic as components deeply rooted in the Entity-Component-System (ECS) pattern. To enhance scalability and resource optimization, we introduce the concept of Ephemeral Rollups (ERs) that overcome the tradeoffs of L2s horizontal scaling. These dedicated runtimes can be customized to provide higher operational speed, configurable ticking mechanisms, provable sessions and gasless transactions without composability-scalability tradeoffs.
@@ 1.1 Game Logic, State, and Transactions
In a fully on-chain game, the game logic resides within smart contracts deployed on the blockchain, ensuring transparency and immutability. It comprises the rules, player interactions, and other essential behaviours that define the gaming experience.
@@
Transactions in this context denote the interactions and exchanges occurring within the game. Players’ actions, asset transfers, and other in-game activities are state transitions recorded on-chain, ensuring their integrity and permanent record. The blockchain’s decentralization makes these interaction transparent, verifiable, and immutable.
@@ 1.2 Outline
In the upcoming sections, 2 and 4, we introduce a performant and scalable SVM-based framework for FOC games and Autonomous Worlds. We dive into some of key features, notably its scalability and the efficient re-use of resources and on-chain components. For an overview of the emerging gaming frameworks and their limitations refer to Appendix A. For an overview of the commonly adopted scalability solutions and their limitations and tradeoffs refer to Appendix B.
@@ 2 Framework Architecture
Developing multiplayer games is a complex endeavour in conventional settings. FOC games further intensify these challenges as we must adhere to registering every interaction on-chain, a limiting factor both in terms of scalability and costs. These limitations primarily arise from its decentralized nature and the consensus mechanisms employed.
Considering a straightforward scenario, a game with 1000 players, with each player’s position updating every 100 ms, results in 10,000 transactions per second. Scaling this to 100,000 players, the rate is leaping to 1,000,000 transactions per second — a volume no blockchain, at the time of writing, can support. Even for a high-speed blockchain like Solana, which might manage such traffic in the future, the demands quickly become insurmountable when considering multiple games and higher numbers of players. From a cost perspective, Solana’s localized fee market does alleviate some concerns by preventing a single game from affecting fee costs for other games (”noisy neighbor” problem). However, even if the individual transaction fees are low, the cumulative cost of such a vast number of transactions can become untenable for a single fee payer (i.e., the game developer).
@@ 2.2 Ephemeral Rollups
The core intuition is that by harnessing the SVM’s account structure and its capacity for parallelization, we can split the app/game state into clusters. Users can lock one or multiple accounts to temporarily transfer the state to an auxiliary layer, which we define as the ”ephemeral rollup”, a configurable dedicated runtime. This process temporarily entrusts the sequencer to modify the accounts in the ephemeral rollup, where the state can be forcefully reverted and unlocked on the L1 if constraints are not satisfied (see section 2.3). Despite this delegation, operations and transactions can still use the delegate accounts as readable on the base layer. Non-delegated accounts remain unaffected and modifiable. The ephemeral rollup operates as a specialized SVM runtime to facilitate transaction processing at an elevated throughput and low-latency. Additionally, this specialized runtime can be customized to include configurations like gasless transactions, quicker block-time, and the inclusion of a ticking mechanism (i.e., an integrated transaction scheduling system like clockwork [clockwork] operated without fees). The entire process is transparent to the end user - a specialized RPC provider can route the transactions to the base layer and the ephemeral rollup(s) in parallel during the game session.
Figure 1 provides an overview of the system. A program defines the instructions and executable logic. Its state consists of various accounts; for a game, PDAs could represent the players’ positions, and another account could be a Chest designated for distributing rewards. The program leverages existing programs on the L1, such as an off-the-shelf leaderboard, a contract to mint NFT or an energy system.
@@
    Monitoring the delegation program, the provisioner detects the provisioning request and promptly launches the corresponding runtime based on the configuration.
@@
    Once the ER concludes (or if the game program forces its closure), account control reverts to the owner program and the provisioner terminates the ER.
Figure 1: The figure showcases the architecture of the Ephemeral Rollups scalability solution, focusing on how a game program manages its state across multiple accounts. The example illustrates two PDAs, each responsible for recording a state, which could be the position of players. These PDAs are then delegated to the Ephemeral Rollup delegation program (DLP). In the background, the Provisioner constantly observes the DLP program, watching for provisioning requests. Upon detection, it dynamically launches an SVM runtime, adapting to specific configuration parameters such as block time and ticking, among others. Within the Ephemeral Rollup (ER) framework, transactions using these delegated accounts are accelerated. After validation, the account state is updated and finalized on the L1.
To illustrate with a practical example, consider Figure 1. A game program might reward players with an NFT reward upon reaching a specified area on a map. PDAs, tracking players’ positions, can be frequently updated within the ER, facilitating a low-latency multiplayer session among players in the ER. When a player reaches the designated map location that triggers a reward, according to the game logic, the client can submit the transaction in the usual manner. If the sequencer doesn’t update rapidly enough, one might first submit an optional state settlement request. Behind the scenes, the Remote Procedure Call (RPC) router relays this transaction to the base layer. Operating on the base layer, the game program can accurately process this transaction and dispense the reward. This is because the position account, which is read-only, reflects the updated location, and modifications can be made to an account which wasn’t delegated through the delegation program.
@@ 2.2.1 Avoiding Fragmentation and Benefits
  • Developers deploy programs to the base layer (e.g., Solana), rather than on a separate chain as it would normally happen with rollups. Programs live on the base layer and can interact with any existing protocol and assets. ERs don’t fragment the existing ecosystem and allow the speed-up of targeted operations without creating an isolated environment.
@@
  • The specialized runtime can accommodate game-specific customizations (e.g., ticking or passive events, typical in games, as opposed to the event-driven runtime of blockchains) without the need to pay gas for fees on those specific accounts.
  • This approach enables a highly scalable system capable of launching rollups on-demand, auto-scaling horizontally to accommodate users performing millions of transactions without the tradeoffs of traditional L2s.
@@ 2.2.2 Reading and Modifying the state
  • If the transaction contains readable accounts from the base/ER layer and all writable accounts are delegated to the ER, the transaction is relayed to the ER RPC.
  • If the transaction includes readable accounts from the base/ER layer and no writable accounts are delegated, the transaction is relayed to the base layer RPC.
@@ 2.3 Security and Computation Trustworthiness
The second point is a common challenge in all rollup architectures, typically addressed through optimistic rollups or ZK rollups. For a detailed description of zk-Rollups, see [ethereum2023zk]; for Optimistic Rollups, see [ethereum2023optimistic]; and for a comparison, see [cryptoslav2021]. [bat] describes ZK channels, a conceptually similar idea for the gaming use-case. As mentioned, transactions history remain available natively in a compressed account on Solana or an external DA layer like Celestia [LazyLedger]). The proposed approach combines these two techniques: it initially performs an optimistic computation for speed, and then produces a ZK proof (e.g., through RISC0 [RISCZero]) for verification. This proof can be efficiently verified on the base layer in a single transaction. Anyone can become a Provisioner for the Ephemeral Rollups sessions by staking a predetermined amount of tokens to the ER program, essentially creating a DePIN infrastructure [solana_depin]. Furthermore, ER participants themselves can act as SVM diet clients [tiny], thereby enhancing the trustworthiness of the network.
@@ 3 SVM as a Multiplayer Game Server
The architecture presented addresses the challenge of processing millions of transactions for an arbitrary number of games. By horizontally scaling the traffic across multiple SVM ERs, it ensures high throughput and low latency, without compromising composability on the base layer and making it suitable for multiplayer gaming scenarios.
@@ 4 ECS, Mapping and Registry
  • Entities: An entity is a general-purpose object typically represented by a unique identifier. It doesn’t directly contain any data or behaviour but serves as an identifier for a collection of components. Entities can be registered in a world instance account on Solana and could themselves be accounts.
  • Components: Raw data structure that represents some aspect or attribute of an entity. For instance, a PositionComponent might just contain x, y, and z coordinates. This is concept is essentially equivalent to accounts on Solana.
  • Systems: Performs the game logic or behaviour by acting upon entities that have specific components. Similar to programs on Solana, which only specify the logic and the accounts they operate on.
Solana’s Program Derived Addresses (PDAs) facilitate the efficient storage and retrieval of components, mimicking a hash table-like structure. [arc] provides an example implementation of an ECS on Solana.
@@ 4.1 Mapping and State listening
The standardized pattern allows for easy integration of game components with a rendering engine to display the game interface. The presented framework can be viewed as an open and permissionless alternative to a backend intended as a multiplayer game server. For visualization and rendering, existing engines such as Unity [unity3d], Unreal Engine [unrealengine], Phaser [phaser], and others can be used. The standardized structure of the components allows for automatic mapping of components and entity properties (abstracting serialization and deserialization, akin to the mechanism of the Anchor framework on Solana [anchor]). State updates can be easily performed and intercepted, providing a mechanism similar to the Observer pattern [observer] to listen to state changes and update the rendering.
@@ 4.2 Public Components Registry
A registry is a complementary infrastructure that works effectively in conjunction with an architecture that leverages the ECS pattern or the general paradigm of Solana. A public registry allows for publishing, discovering, and reusing pieces of logic (such as systems and data structures) shared across apps and games. Every program deployed on Solana can be perceived as a system that delineates the logic and receives accounts as input for computation. Specifically for apps and games, we envision a registry housing typical developmental logic - a collision system, an energy system, a movement system etc. — commonly found in traditional engines. Thus, the registry serves as a single hub where developers can discover, publish, and utilize systems and components. This facilitates the swift assembly of games by developers without the need to rewrite the entire underlying logic from scratch.
@@ 4.3 Game Loops, Ticking and FRP
  • Update the game state (e.g., physics, AI, object positions).
A client-side engine would also include the rendering step. As mentioned in section 2.2, continuous state updates can be achieved using a transaction scheduling mechanism like Clockwork. While it is a viable solution, it presents some issues such as the cost of fees and the uncertainty of the exact block/timestamp of the transaction execution. These issues are much less prominent in Ephemeral Rollup, where the cost of state update transactions can be lifted, and the block time can be configured to be in the range of 10 to 50ms.
While having game loops constitutes the most adopted solution for most game engines, there are also less explored alternatives. There are Haskell engines that instead leverage the functional programming paradigm, such as Yampa [Cheong2005FunctionalPA] and implementations of the functional and asynchronous ECS pattern, such as Apecs [apecs]. The underlying idea is that values are derived through a function instead of evolving the state based on update loops (which can depend on input, time, or other sources). A simple example is an energy system that depends on time, e.g., a function that increases energy by one point every 30 seconds. This paradigm aligns perfectly with the realm of blockchain, where transactions are inherently asynchronous and where the game instance potentially lives forever on the blockchain, evolving over time. The idea is conceptually identical to Functional Reactive Programming (FRP) and Functional Reactive Animation [elliott1997functional], which is a programming paradigm that deals with asynchronous data streams and changes in data over time.
Although it is theoretically possible to write an engine that fully utilizes this paradigm, avoiding an update loop, in practice, a combination of the two techniques allows for greater configurability and the ability to use the suitable paradigm considering the game mechanics being developed. For example, it is possible to update the position of one non-performing character (NPC) with a loop and derive the position of $n$ other NPCs. The framework will evolve supporting both paradigms.
@@ 5 Conclusion
In this article, we have presented the design and architecture of Ephemeral Rollups, a performant and scalable solution for fully on-chain games and Autonomous Worlds. The framework adopts an Entity Component System (ECS) architecture to facilitate the creation and reuse of on-chain components and logic. By leveraging the unique features of the Solana Virtual Machine — its account structure, the delegation of account modifications, and parallelization capabilities - the framework proposes Ephemeral Rollups as a promising path to scale FOCG. This design enables fast, specialized runtimes that auto-scale horizontally to meet demand without creating isolated environments. As a result, the system maintains composability with existing protocols, tooling and runtime advancements at the base layer without any state fragmentation. It’s worth noting that while the solution is designed for fully on-chain games, this mechanism can be broadly applied to various other use cases.
@@ Appendix A Emerging gaming frameworks
Recently, several frameworks have emerged in an attempt to simplify the development of this new domain of games. Each of these ”on-chain game engines” possesses distinct features and iterates along different verticals to facilitate the creation and management of FOC games, with various degrees of trade-offs. This section delineates the technical nuances and provides a comparative analysis of these frameworks.

### v4

@@ Abstract
In the realm of open and composable gaming, we envision platforms where users actively expand, create, engage, and immerse themselves in a rich world of entertainment. One promising avenue for achieving this vision is through fully on-chain (FOC) games, where both game state and logic reside on the blockchain, maximizing composability. However, we must grapple with inherent limitations and trade-offs, particularly in terms of costs and scalability. This paper proposes a framework that leverages the Solana Virtual Machine (SVM) to scale FOC games without state fragmentation or compromised trust assumptions. The framework introduces a systematic approach for discovering, utilizing, and publishing modular pieces of logic as components deeply rooted in the Entity-Component-System (ECS) pattern. To enhance scalability and resource optimization, we introduce the concept of Ephemeral Rollups (ERs) that overcome the tradeoffs of L2 horizontal scaling. These dedicated runtimes can be customized to provide higher operational speed, configurable ticking mechanisms, provable sessions and gasless transactions without composability-scalability tradeoffs.
@@ 1.1 Game Logic, State, and Transactions
In a fully on-chain game, the game logic resides within smart contracts deployed on the blockchain, ensuring transparency and immutability. It comprises the rules, player interactions, and other behaviors that define the gaming experience.
@@
Transactions in this context denote the interactions and exchanges that occur within the game. Players’ actions, asset transfers, and other in-game activities are state transitions recorded on-chain, ensuring their integrity and permanent record. The blockchain’s decentralization makes these interaction transparent, verifiable, and immutable.
@@ 1.2 Outline
In the upcoming sections, 2 and 4, we introduce a performant and scalable SVM-based framework for FOC games and Autonomous Worlds. We dive into some of key features, notably its scalability and the efficient reuse of resources and on-chain components. For an overview of the emerging gaming frameworks and their limitations, refer to Appendix A. For an overview of the commonly adopted scalability solutions and their limitations and trade-offs, see the Appendix B.
@@ 2 Framework Architecture
Developing multiplayer games is a complex endeavour in conventional settings. FOC games further intensify these challenges, as we must adhere to registering every interaction on-chain, a limiting factor both in terms of scalability and costs. These limitations primarily arise from its decentralized nature and the consensus mechanisms employed.
Considering a straightforward scenario, a game with 1000 players, with each player’s position updating every 100 ms, results in 10,000 transactions per second. Scaling this to 100,000 players, the rate leaps to 1,000,000 transactions per second — a volume no blockchain, at the time of writing, can support. Even for a high-speed blockchain like Solana, which might manage such traffic in the future, the demands quickly become insurmountable when considering multiple games and higher numbers of players. From a cost perspective, Solana’s localized fee market does alleviate some concerns by preventing a single game from affecting fee costs for other games (”noisy neighbor” problem). However, even if the individual transaction fees are low, the cumulative cost of such a vast number of transactions can become untenable for a single fee payer (i.e., the game developer).
@@ 2.2 Ephemeral Rollups
The core intuition is that by harnessing the SVM’s account structure and its capacity for parallelization, we can split the app/game state into clusters. Users can lock one or multiple accounts to temporarily transfer the state to an auxiliary layer, which we define as the ”ephemeral rollup”, a configurable dedicated runtime. This process temporarily allows the sequencer to modify accounts within the ephemeral rollup, with the state being forcibly reverted and unlocked on the L1 if constraints are not met (see section 2.3). Despite this delegation, operations and transactions can still use the delegate accounts as readable on the base layer. Non-delegated accounts remain unaffected and modifiable. The ephemeral rollup operates as a specialized SVM runtime to facilitate high-throughput, low-latency transaction processing. Additionally, this specialized runtime can be customized to include configurations like gasless transactions, quicker block-time, and the inclusion of a ticking mechanism (i.e., an integrated transaction scheduling system like clockwork [clockwork] operated without fees). The entire process is transparent to the end user - a specialized RPC provider can route the transactions to the base layer and the ephemeral rollup(s) in parallel during the game session.
Figure 1 provides an overview of the system. A program defines executable logic, with its state consisting of various accounts; for a game, PDAs could represent the players’ positions, and another account could be a Chest designated for distributing rewards. The program leverages existing programs on the L1, such as an off-the-shelf leaderboard, a contract to mint NFT or an energy system.
@@
    Monitoring the delegation program, the provisioning request, and immediately launches the corresponding runtime based on the configuration.
@@
    Once the ER concludes (or if the game program forces closure), account control reverts to the owner program and the provisioner terminates the ER.
Figure 1: The figure showcases the architecture of the Ephemeral Rollups scalability solution, focusing on how a game program manages its state across multiple accounts. The example illustrates two PDAs, each responsible for recording a state, which could be the position of players. These PDAs are then delegated to the Ephemeral Rollup delegation program (DLP). In the background, the Provisioner constantly observes the DLP program, watching for provisioning requests. Upon detection, it dynamically launches an SVM runtime, adjusting to specific configuration parameters like block time, ticking, and others. Within the Ephemeral Rollup (ER) framework, transactions using these delegated accounts are accelerated. After validation, the account state is updated and finalized on the L1.
To illustrate with a practical example, consider Figure 1. A game program might reward players with an NFT reward upon reaching a specified area on a map. PDAs, tracking players’ positions, can be frequently updated within the ER, facilitating a low-latency multiplayer session among players in the ER. When a player reaches the designated map location that triggers a reward, according to the game logic, the client can submit the transaction in the usual manner. If the sequencer doesn’t update rapidly enough, users may first submit an optional state settlement request. Behind the scenes, the Remote Procedure Call (RPC) router relays this transaction to the base layer. Operating on the base layer, the game program can accurately process this transaction and dispense the reward. This is because the position account, which is read-only, reflects the updated location, and modifications can be made to an account that wasn’t delegated through the delegation program.
@@ 2.2.1 Avoiding Fragmentation and Benefits
  • Developers deploy programs to the base layer (e.g., Solana), rather than on a separate chain as it would normally happen with rollups. Programs reside on the base layer and can interact with any existing protocol and assets. ERs don’t fragment the existing ecosystem and allow the speed-up of targeted operations without creating an isolated environment.
@@
  • The specialized runtime can accommodate game-specific customizations (e.g., ticking or passive events, typical in games, as opposed to the event-driven runtime of blockchains) without the need for paying gas fees.
  • This approach enables a highly scalable system capable of launching rollups on-demand and auto-scaling horizontally to handle millions of transactions without the tradeoffs of traditional L2s.
@@ 2.2.2 Reading and Modifying the state
  • If the transaction contains readable accounts from the base or ER layer and all writable accounts are delegated to the ER, the transaction is relayed to the ER RPC.
  • If the transaction includes readable accounts from the base or ER layer and no writable accounts are delegated, the transaction is relayed to the base layer RPC.
@@ 2.3 Security and Computation Trustworthiness
The second point is a common challenge in all rollup architectures, typically addressed through optimistic rollups or ZK rollups. For a detailed description of zk-Rollups, see [ethereum2023zk]; for Optimistic Rollups, see [ethereum2023optimistic]; and for a comparison, see [cryptoslav2021]. [bat] describes ZK channels, a conceptually similar idea for the gaming use-case. As mentioned, transaction history remains available natively in a compressed account on Solana or an external DA layer like Celestia [LazyLedger]). The proposed approach combines these two techniques: it initially performs an optimistic computation for speed, and then produces a ZK proof (e.g., through RISC0 [RISCZero]) for verification. This proof can be efficiently verified on the base layer in a single transaction. Anyone can become a Provisioner for the Ephemeral Rollups sessions by staking a predetermined amount of tokens to the ER program, essentially creating a DePIN infrastructure [solana_depin]. Furthermore, ER participants themselves can act as SVM diet clients [tiny], thereby enhancing the trustworthiness of the network.
@@ 3 SVM as a Multiplayer Game Server
The architecture presented addresses the challenge of processing millions of transactions for an arbitrary number of games. By horizontally scaling the traffic across multiple SVM ERs, it ensures high throughput and low latency, without compromising composability on the base layer, making it ideal for multiplayer gaming scenarios.
@@ 4 ECS, Mapping and Registry
  • Entities: An entity is a general-purpose object typically represented by a unique identifier. It does not directly contain any data or behavior but serves as an identifier for a collection of components. Entities can be registered within a world instance account on Solana and could also be represented as individual accounts.
  • Components: Raw data structure that represents some aspect or attribute of an entity. For example, a PositionComponent might just contain x, y, and z coordinates. This concept is similar to how accounts function on Solana.
  • Systems: Perform the game logic or behavior by acting upon entities with specific components, much like how programs on Solana define the logic and operate on specific accounts.
Solana’s Program Derived Addresses (PDAs) facilitate the efficient storage and retrieval of components, mimicking the structure of a hash table. [arc] provides an example implementation of an ECS on Solana.
@@ 4.1 Mapping and State listening
The standardized pattern allows for easy integration of game components with a rendering engine to display the game interface. The presented framework can be viewed as an open and permissionless alternative to a backend intended as a multiplayer game server. For visualization and rendering, existing engines such as Unity [unity3d], Unreal Engine [unrealengine], Phaser [phaser], and others can be used. The standardized structure of the components allows for automatic mapping of components and entity properties (abstracting serialization and deserialization, akin to the mechanism of the Anchor framework on Solana [anchor]). State updates can be easily executed and monitored, providing a mechanism similar to the Observer pattern [observer] to listen to state changes and update the rendering.
@@ 4.2 Public Components Registry
A registry is a complementary infrastructure that works effectively in conjunction with an architecture that leverages the ECS pattern or the general Solana paradigm. A public registry allows for publishing, discovering, and reusing pieces of logic (such as systems and data structures) shared across apps and games. Every program deployed on Solana can be perceived as a system that delineates the logic and receives accounts as input for computation. Specifically for apps and games, we envision a registry housing typical developmental logic - a collision system, an energy system, a movement system etc. — commonly found in traditional engines. Thus, the registry serves as a hub where developers can discover, publish, and utilize systems and components. This facilitates the swift assembly of games by developers without the need to rewrite the entire underlying logic from scratch.
@@ 4.3 Game Loops, Ticking and FRP
  • Process game state updates (e.g., physics, AI, object positions).
A client-side engine would also include the rendering step. As mentioned in section 2.2, continuous state updates can be achieved using a transaction scheduling mechanism such as Clockwork. While it is a viable solution, it presents some issues such as the cost of fees and the uncertainty of the exact block/timestamp of the transaction execution. These issues are less prominent in Ephemeral Rollups, where state update transaction costs can be eliminated, and block time can be configured between 1 and 50ms.
While having game loops constitutes the most adopted solution for most game engines, there are also less explored alternatives. There are Haskell engines that instead leverage the functional programming paradigm, such as Yampa [Cheong2005FunctionalPA] and implementations of the functional and asynchronous ECS pattern, such as Apecs [apecs]. The underlying idea is that values are derived through a function instead of evolving the state based on update loops (which can depend on input, time, or other sources). A simple example is an energy system that depends on time, e.g., a function that increases energy by one point every 30 seconds. This paradigm aligns perfectly with the realm of blockchain, where transactions are inherently asynchronous and where the game instance potentially lives forever on the blockchain, evolving over time. This is conceptually similar to Functional Reactive Programming (FRP) and Functional Reactive Animation [elliott1997functional], which is a programming paradigm that deals with asynchronous data streams and changes in data over time.
Although it is theoretically possible to write an engine that fully utilizes this paradigm, avoiding an update loop, in practice, a combination of the two techniques allows for greater configurability and the ability to use the suitable paradigm considering the game mechanics being developed. For example, it is possible to update the position of one non-performing character (NPC) with a loop and derive the position of $n$ other NPCs. The framework will evolve to support both paradigms.
@@ 5 Conclusion
In this article, we have presented the design and architecture of Ephemeral Rollups, a performant and scalable solution for fully on-chain games and Autonomous Worlds. The framework adopts an Entity Component System (ECS) architecture to facilitate the creation and reuse of on-chain components and logic. By leveraging the unique features of the Solana Virtual Machine — its account structure, the delegation of account modifications, and parallelization capabilities — the framework proposes Ephemeral Rollups as a promising path to scale FOC Games. This design enables fast, specialized runtimes that auto-scale horizontally to meet demand without creating isolated environments. As a result, the system maintains composability with existing protocols, tooling and runtime advancements at the base layer without any state fragmentation. While this solution is designed for fully on-chain games, it’s important to note that the mechanism can be broadly applied to various other use cases.
@@ Appendix A Emerging gaming frameworks
Recently, several frameworks have emerged in an attempt to simplify the development of this new domain of games. Each of these ’on-chain game engines’ has distinct features and approaches to facilitate the creation and management of FOC games, each with varying trade-offs. This section delineates the technical nuances and provides a comparative analysis of these frameworks.